[ An open box is constructed from a rectangular sheet of cardboard that measures 12 inches by
18 inches. Square pieces are cut out of the corners so that when the box is formed it will be 3/4
inches deep. What is the volume of the box? Express your answer as a mixed number. ]

Answer by TimothyLamb(4379)

size of box bottom:
w = 12 - 2(3/4)
w = (48/4) - (6/4)
w = (42/4)
w = (21/2)
---
L = 18 - 2(3/4)
L = (66/4)
L = (33/2)
---
v = Lwh
v = (33/2)(21/2)(3/4)
v = (2079/16)
v = (2079/16)
v = 129.9375
---
answer:
v = 129 + 15/16
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php